Suppose a wave function were assumed to be given by ψ = A from x = a to x = b, ψ = 0 elsewhere. Normalization yields what value for A?

1

b - a

(b - a)-1

(b - a)-1/2
Answer #1

Wavefunction = A

Normalisation

dx = 1

[ b-a] = 1

=1/(b-a)

Taking square root on both sides

A= 1/(b-a)^1/2

A=( b-a)^-1/2

Last option is correct